OpenStudy (raheelafzaal):

|-4| + |4| - 4 + 4 = ?

OpenStudy (doc.brown):

Do the lines mean absolute value?

OpenStudy (raheelafzaal):

|-4| + |4| - 4 + 4 = ? 0 2 4 8

OpenStudy (anonymous):

4 + 4 - 4 + 4 = 8 ABS value bars remove negatives

OpenStudy (raheelafzaal):

only negatives ll remove?

OpenStudy (doc.brown):

I like doing it like this, it's more fun:\[4+4-4+4\]\[4+4\cancel{-4}\cancel{+4}\]\[4+4=8\]

OpenStudy (anonymous):

Absolute value is thinking in terms of magnitude. If I drew arrows from 0 to both -4 and 4 on a number line, each arrow has a magnitude of 4, but one has a positive direction (to the right) and the other negative (to the left). So, absolute value is just asking for magnitude. 

OpenStudy (anonymous):

therefore \[|-4| = |4| = 4\]

OpenStudy (raheelafzaal):

@mathbrz thanks now i understand it

OpenStudy (anonymous):

yw glad to help :)